OpenStudy (anonymous):

Find the arc length of the curve over the given interval. (Round your answer to three decimal places.) y = ln x, [1, 3]

OpenStudy (anonymous):

first find \(f'(x)\)

OpenStudy (anonymous):

let me know when you have that

OpenStudy (anonymous):

1/x

OpenStudy (anonymous):

ok good now square it

OpenStudy (anonymous):

1/x^2

OpenStudy (anonymous):

ok, now add 1

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

this is the easy part, the hard part is coming up

OpenStudy (anonymous):

did you get \[\frac{x^2+1}{x^2}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok and the square root of that is \[\frac{\sqrt{x^2+1}}{x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now we gotta figure out \[\int_1^3\frac{\sqrt{x^2+1}}{x}dx\]

OpenStudy (anonymous):

2.3020

OpenStudy (anonymous):

at this point i would just cheat and use a calculator to get the answer, finding the anti derivative is going to be a pain, require a substitution \[u=\tan(x)\] to get rid of the radical

OpenStudy (anonymous):

that is what i get too

OpenStudy (anonymous):

Okay! Thank you!

OpenStudy (anonymous):

yw